OpenStudy (anonymous):

The first prize in a lottery is $100 000. Each succeeding winning number pays 40% as much as the winning number before. How much is paid out in prizes if 6 numbers are drawn?

OpenStudy (anonymous):

100,000*0.40=40,000(1st), 40,000*0.40=16,000(2nd), 16,000*0.40=6,400(3rd), 6,400*0.40=2,560(4th), 2,560*0.40=1,024(5th), 1,024*0.40=409.60(6th) hope this helps:)

OpenStudy (anonymous):

Would there be a formula? Cause we're dealing with Geometric Series right now :(

OpenStudy (anonymous):

I would probably have to say that it would be the amount*your percentage to get an answer like that. Please bear with me I'm just getting used to this program:)

OpenStudy (anonymous):

Its okay, thank you for the help though! :)

OpenStudy (anonymous):

Does this make any sense to u???

OpenStudy (anonymous):

I was just asking if I were to use the Sn = a(r^n - a)/r-1

OpenStudy (anonymous):

I'm sorry I havn't done that before and I'm actually in algebra right now so its hard to think geometry. I have to do my homework now sorry.

OpenStudy (anonymous):

Its alright, thanks for your help :)